40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 8 20:48:06 2017, gautam, Update, SUS, ITMX stuck 
    Reply  Thu Aug 10 09:31:57 2017, Steve, Update, SUS, ITMX sensor voltage 9daysITMX.pngvacGlitchITMX.png
Message ID: 13173     Entry time: Tue Aug 8 20:48:06 2017     Reply to this: 13182
Author: gautam 
Type: Update 
Category: SUS 
Subject: ITMX stuck 

Somewhere between CDS model restarts and the IFO venting, ITMX got stuck.

I shook it loose using the usual bias slider technique. It appears to be free now, I was able to lock the green beam on a TEM00 mode without touching the green input pointing. The ITMX Oplev spot has also returned to within its MEDM display bounds.

ELOG V3.1.3-